# : Timing Interruption Monitoring and Security Assessment for Fronthaul Environments

### v1

@@ 7.1 Heuristic Detection Results
Testing under the same conditions we use in Section 9.2 showed that while the heuristic effectively detects Announce attacks, it struggles with overall accuracy. The primary issue is its sensitivity to individual packets: legitimate packets observed during an attack can reset the heuristic, causing it to miss subsequent malicious packets. Results, as shown in Fig. 8, indicate poor overall performance due to these limitations.
8 ML-Based Detection
@@
8.1 ML Model Selection
We used our digital twin environment to generate training data, as described in Sec. 6.2. The data was split into equal-sized chunks (1000 messages each) without overlaps. We randomly selected 80% of the chunks for training, 10% for validation, and 10% for testing. Upon acceptance, we will open-source our complete data set and training pipeline.
We initially evaluated a wide range of ML models, as shown in Table III, using the same test traces for all offline accuracy tests. All ML models tested had higher accuracy than the rule-based approach in Sec. 7.1. However, accuracy alone is not sufficient; the impact of False Negatives (FNs) (or missing an attack) is catastrophic, as shown in Section 6.1, while the impact of False Positives (FPs) is much lower. Therefore, we also examine each model’s recall. Given the clear performance advantage of the CNN and Transformer models, we selected both for further evaluation. The complete confusion matrices for the initial offline results for these models are shown in Fig. 9.
@@ 8.1 ML Model Selection
8.2 ML Model Description
Here we briefly describe the design of the two best performing models, the CNN and transformer. A complete description of each model along with code used will be made open source upon acceptance of this article. All the models use the same six input features, : Ethernet source and destination addresses; packet size; PTP sequence ID; PTP message type; and inter-arrival time. The MAC address helps identify flows and packet direction. The PTP message type and sequence ID are crucial for understanding the nature and order of PTP messages, aiding in the detection of deviations from expected behavior. Packet length can indicate malformed or malicious packets with extra payload. Inter-arrival time, specified by the O-RAN ALLIANCE for both announce and sync messages, is another key indicator of potential attacks, though minor variations between packets are expected.
@@ 8.2 ML Model Description
Our transformer-based model (see Fig. 10) uses only the encoder portion of a full transformer architecture. It processes input sequences with fixed lengths, varied during training and evaluation to include lengths of packets. These lengths balance the trade-off between incorporating more historical data (and requiring more memory) and improving accuracy and recall. We also treated the number of attention heads as a hyper-parameter with two configurations. The transformer encoder outputs a feature matrix of dimension , which is then fed into a fully connected linear layer with a ReLU activation function, followed by a dropout layer with a probability of 0.2 and a final linear layer with a sigmoid activation function. The model produces a single binary classification for the sequence.
@@
For both the CNN and transformer models, we employed a custom Weighted Binary Cross-Entropy Loss function to address class imbalance and unequal error costs. However, since the initial setting of 1 for both weights yielded satisfactory results, we did not adjust these weights further. We used the Adam optimizer with an initial learning rate of 0.001 and configured a StepLR scheduler to reduce the learning rate by a factor of 0.1 every 10 epochs. Early stopping was enabled, with monitoring for validation loss improvement and a patience parameter set to 20 epochs.
8.3 Deploying
To test our proposed solution, we built a detection pipeline with three parallel components: packet acquisition and filtering, pre-processing, and decision-making. The pipeline captures Ethernet frames, filters for PTP packets, and extracts key features, such as MAC addresses, length, sequence ID, message type, and inter-arrival timestamps. The pre-processing stage maps observed MAC addresses to integers, enabling the model to learn traffic patterns rather than specific addresses. The decision-making block uses a sliding window approach to classify each sequence as benign or malicious, using either an ML model or a rule-based algorithm. We use a stride of two for our sliding window, so at most two malicious packets are sent before our model detects the attack.
9 Detection Results
@@
9.1 Production-ready Detection Accuracy
@@
We evaluated the detection pipeline using a trace (.pcap) captured during a successful attack in the production-ready environment, as illustrated in Fig. 5. The results demonstrate that the transformer model with 2 attention heads and a slice length of 32 significantly outperformed other transformer configurations, achieving an accuracy exceeding 99%. The confusion matrix in Fig. 11(a) illustrates the model’s strong performance, highlighting its ability to accurately distinguish between malicious and benign traffic in a real-world setting.
@@
In contrast, Fig. 11(b) shows that the performance of the CNN model was substantially degraded when applied to the new environment. This suggests that the CNN struggled to generalize to unseen data and failed to adapt to the subtle contextual variations present in the production environment. The transformer model, however, maintained its robustness, effectively identifying the nuanced patterns and indicators of the attack, even in a more complex and dynamic setting. This reinforces the transformer’s ability to generalize well to new conditions, capturing intricate relationships across sequences that may vary from one environment to another.
9.2 Digital Twin Detection Accuracy
@@
To further validate our ML model, we deployed in a digital twin environment that simulates the production-ready network. This setup provides a realistic testbed for evaluating the model’s performance under a wide variety of conditions. While running full PTP with ptp4l, we keep system time updates on the DU and attacker disabled, allowing us to log prediction outcomes with accurate DU timestamps. The attacker logs the attack type and its timestamps, enabling us to measure decision accuracy. Analysis reveals that we can detect attacks with just one or two malicious PTP packets reaching the DU.
@@
During the experiment, several tests were performed. At the start of each test, the DU, RU, and attacker exchanged initial messages to confirm readiness. Once prepared, they initiated a 5-minute test. Throughout the test, the DU continuously logged predictions with timestamps, while the attacker recorded the attack types, along with their start and end times, and the duration of both the attack and recovery phases. This cycle repeated for the entire test duration, alternating between attack and recovery phases. Recall that during training, we used cycles of 30/30, 40/20, 50/10 seconds, and continuous attack, as described in Section 6.2. To evaluate the models’ ability to generalize to altered attack patterns, the attacker in this experiment randomly selected an attack duration from the continuous range seconds and a recovery duration from the continuous range seconds. The detection results from this experiment are shown in Fig. 12. The transformer demonstrated excellent resilience against the reshaped attack, while the CNN completely failed to detect it. This outcome highlights the transformer’s robust ability to adapt to previously unseen attack strategies, confirming greater suitability for real-world deployment where attack patterns can be unpredictable.
10 Discussion
@@
The Open RAN framework’s push towards increased virtualization and softwarization [51, 52] necessitates security solutions that align with this trend. Prong D (Sec. 2.3), which involves monitoring, is well-suited to the softwarized DU environment. However, future work should prioritize several additional key areas.
@@
Attacks in Other Synchronization Topologies. Our focus on LLS-C2 and LLS-C3, due to their high vulnerability, does not encompass all risks. For instance, jamming GNSS signals could affect LLS-C4, and supply chain attacks on primary reference clocks pose additional threats. Network redundancy, as suggested in Prong C, enhances protection and resilience, potentially mitigating such threats through redundant paths and clocks [53, 54, 16, 36].
@@
Additional Attacks Against PTP. While currently addresses two types of attacks, other threats like packet removal and selective delay manipulation also merit investigation [15]. Expanding to include multi-class classification for various attack types could refine threat detection and response strategies.
@@
11 Conclusion

### v2

@@ 7.1. Heuristic Detection Results
Testing under the same conditions we use in SectionΒ 9.2 showed that while the heuristic effectively detects Announce attacks, it struggles with overall accuracy. The primary issue is its sensitivity to individual packets: legitimate packets observed during an attack can reset the heuristic, causing it to miss subsequent malicious packets. Results, as shown in Fig.Β 8, indicate poor overall performance due to these limitations.
8. ML-Based Detection
@@
8.1. ML Model Selection
We used our digital twin environment to generate training data, as described in Sec.Β 6.2. The data was split into equal-sized chunks (1000 messages each) without overlaps. We randomly selected 80% of the chunks for training, 10% for validation, and 10% for testing. Upon acceptance, we will open-source our complete data set and training pipeline.
We initially evaluated a wide range of ML models, as shown in TableΒ 3, using the same test traces for all offline accuracy tests. All ML models tested had higher accuracy than the rule-based approach in Sec.Β 7.1. However, accuracy alone is not sufficient; the impact of False Negatives (FNs) (or missing an attack) is catastrophic, as shown in SectionΒ 6.1, while the impact of False Positives (FPs) is much lower. Therefore, we also examine each modelβs recall.
We do recognize that in operational environments, even a modest false positive rate can generate frequent alerts and contribute to operator fatigue, especially when attacks are rare. Heuristic methods may have an advantage of allowing direct adjustment of detection thresholds to accommodate different tolerance for false positives. Although our ML-based approach does not expose simple rule thresholds, the effective sensitivity of the model can be tuned through several mechanisms, such as adjusting the decision threshold on the output probabilities, retraining with modified class weights, or employing cost-sensitive loss functions. These mechanisms allow operators to adapt the false positive/false negative tradeoff to specific deployment needs while maintaining the modelβs ability to learn complex, non-linear behaviors that heuristic rules cannot capture.
Given the clear performance advantage of the CNN and Transformer models, we selected both for further evaluation. The complete confusion matrices for the initial offline results for these models are shown in Fig.Β 9. Additional details about all models are provided in AppendixΒ A.3.
@@ 8.1. ML Model Selection
8.2. ML Model Description
Here we briefly describe the design of the two best performing models, the CNN and transformer. A complete description of each model (see A.3) along with code used will be made open source upon acceptance of this article. All the models use the same six input features, : Ethernet source and destination addresses; packet size; PTP sequence ID; PTP message type; and inter-arrival time. The MAC address helps identify flows and packet direction. The PTP message type and sequence ID are crucial for understanding the nature and order of PTP messages, aiding in the detection of deviations from expected behavior. Packet length can indicate malformed or malicious packets with extra payload. Inter-arrival time, specified by the O-RAN ALLIANCE for both announce and sync messages, is another key indicator of potential attacks, though minor variations between packets are expected.
@@ 8.2. ML Model Description
Our transformer-based model (see Fig.Β 10) uses only the encoder portion of a full transformer architecture. It processes input sequences with fixed lengths, varied during training and evaluation to include lengths of packets. These lengths balance the trade-off between incorporating more historical data (and requiring more memory) and improving accuracy and recall. We also treated the number of attention heads as a hyper-parameter with two configurations. The transformer encoder outputs a feature matrix of dimension , which is then fed into a fully connected linear layer with a ReLU activation function, followed by a dropout layer with a probability of 0.2 and a final linear layer with a sigmoid activation function. The model produces a single binary classification for the sequence.
@@
Each sequence of length (used for both CNN and Transformer models) was classified as malicious if any packet in the sequence was labeled as malicious, allowing the models to learn patterns of attack behavior over time. For both the CNN and transformer models, we employed a custom Weighted Binary Cross-Entropy Loss function to address class imbalance and unequal error costs. However, since the initial setting of 1 for both weights yielded satisfactory results, we did not adjust these weights further. In practice, these weights, which correspond to the decision threshold, could be modified to adjust the sensitivity of detection. For example, deployments with more stable network baselines may prefer higher precision to reduce false alarms, whereas high-security environments might prioritize recall to minimize missed attacks. Exploring adaptive threshold tuning or online re-weighting in future work could further improve deployability across diverse scenarios. We used the Adam optimizer with an initial learning rate of 0.001 and configured a StepLR scheduler to reduce the learning rate by a factor of 0.1 every 10 epochs. Early stopping was enabled, with monitoring for validation loss improvement and a patience parameter set to 20 epochs.
Training time depends on both the size of the dataset and the hardware used. However, training is performed offline, making it a one-time cost that does not impact real-time performance. On modern hardwareβsuch as a single NVIDIA A100 GPU on a DGX machineβtraining never exceeded 30 minutes. This cost is negligible in practice, especially when amortized across many long-term deployments. In the broader context, it is a negligible cost to pay for adding robust security to production environments.
8.3. Deploying
We deploy as a passive detection pipeline to monitor PTP synchronization traffic in the fronthaul segment of the O-RAN architecture. It can run directly on the DU, observing the Ethernet interface connected to the fronthaul switch, or on a separate device connected to a mirror port in the switched networkβparticularly valuable in LLS-C2 and LLS-C3 profiles where synchronization messages traverse multiple hops and are most vulnerable to attack.
The pipeline consists of three modular stages: packet acquisition and filtering, pre-processing, and decision-making, as shown in Fig.Β 11. The first stage captures Ethernet frames and filters for PTP packets, extracting six features: source and destination MAC addresses, frame length, PTP sequence ID, message type, and inter-arrival time. These are passed via thread-safe queues to the pre-processing stage, which maps MAC addresses to integers to abstract away device-specific identifiers and enable generalizable traffic pattern learning.
The decision-making stage maintains a sliding window of size and evaluates each sequence using either a machine learning model or a rule-based algorithm. After every two new PTP packets, the window advances, and a new classification decision is made. The modelβs detection process is designed to operate in real time, with low-latency inference on both the CNN and Transformer modelsβeven on moderate hardware. While the Transformer architecture is computationally heavier due to its attention mechanism, both models sustain real-time decision-making without introducing bottlenecks in both of our testbeds. This approach enables prompt detection of anomalies while keeping latency low. The modular design also allows each stage to be updated independently, supporting flexible deployment and future model improvements.
9. Detection Results
@@
9.1. Production-ready Detection Accuracy
@@
We evaluated the detection pipeline using a trace (.pcap) captured during a successful attack in the production-ready environment, as illustrated in Fig.Β 5. The results demonstrate that the transformer model with 2 attention heads and a slice length of 32 significantly outperformed other transformer configurations, achieving an accuracy exceeding 99%. The confusion matrix in Fig.Β 12(a) illustrates the modelβs strong performance, highlighting its ability to accurately distinguish between malicious and benign traffic in a real-world setting.
@@
In contrast, Fig.Β 12(b) shows that the performance of the CNN model was substantially degraded when applied to the new environment. This suggests that the CNN struggled to generalize to unseen data and failed to adapt to the subtle contextual variations present in the production environment. The transformer model, however, maintained its robustness, effectively identifying the nuanced patterns and indicators of the attack, even in a more complex and dynamic setting. This reinforces the transformerβs ability to generalize well to new conditions, capturing intricate relationships across sequences that may vary from one environment to another.
While the CNN exhibited slightly lower false positives in controlled digital twin experiments, its inability to generalize to the production environment led to a significant drop in overall detection reliability. In contrast, the transformerβs small increase in false positives is outweighed by its robustness to environmental shifts and unseen attack patterns. In practical deployments where detection failures are more costly than occasional false alarms, this tradeoff favors the transformer as the more stable and effective choice.
9.2. Digital Twin Detection Accuracy
@@
To further validate our ML model, we deployed in a digital twin environment that simulates the production-ready network. This setup provides a realistic testbed for evaluating the modelβs performance under a wide variety of conditions. While running full PTP with ptp4l, we keep system time updates on the DU and attacker disabled, allowing us to log prediction outcomes with accurate DU timestamps. The attacker logs the attack type and its timestamps, enabling us to measure decision accuracy. Analysis reveals that we can detect attacks with just one or two malicious PTP packets reaching the DU.
@@
During the experiment, several tests were performed. At the start of each test, the DU, RU, and attacker exchanged initial messages to confirm readiness. Once prepared, they initiated a 5-minute test. Throughout the test, the DU continuously logged predictions with timestamps, while the attacker recorded the attack types, along with their start and end times, and the duration of both the attack and recovery phases. This cycle repeated for the entire test duration, alternating between attack and recovery phases. Recall that during training, we used cycles of 30/30, 40/20, 50/10 seconds, and continuous attack, as described in SectionΒ 6.2. To evaluate the modelsβ ability to generalize to altered attack patterns, the attacker in this experiment randomly selected an attack duration from the continuous range seconds and a recovery duration from the continuous range seconds. The detection results from this experiment are shown in Fig.Β 13. The transformer demonstrated excellent resilience against the reshaped attack, while the CNN completely failed to detect it. This outcome highlights the transformerβs robust ability to adapt to previously unseen attack strategies, confirming greater suitability for real-world deployment where attack patterns can be unpredictable.
10. Discussion
@@
The Open RAN frameworkβs push towards increased virtualization and softwarizationΒ (afolabi2018network; ko2024edgeric) necessitates security solutions that align with this trend. Prong D (Sec.Β 2.3), which involves monitoring, is well-suited to the softwarized DU environment. However, future work should prioritize several additional key areas.
@@
Attacks in Other Synchronization Topologies. Our focus on LLS-C2 and LLS-C3, due to their high vulnerability, does not encompass all risks. For instance, jamming GNSS signals could affect LLS-C4, and supply chain attacks on primary reference clocks pose additional threats. Network redundancy, as suggested in Prong C, enhances protection and resilience, potentially mitigating such threats through redundant paths and clocksΒ (xing2023enabling; lazarev2023resilient; shi2023ms; finkenzeller2024ptpsec).
@@
Additional Attacks Against PTP. While currently addresses two types of attacks, other threats like packet removal and selective delay manipulation also merit investigationΒ (s_plane_survey). Expanding to include multi-class classification for various attack types could refine threat detection and response strategies.
@@
Our detection-based approach is designed to complement these preventive measures. Unlike cryptographic or hardware-based defenses, which may require system upgrades or protocol changes, operates on traffic patterns alone and can be deployed as a lightweight, software-only monitor in existing networks. While authentication mechanisms aim to prevent certain attacks, helps identify when those defenses may be necessary, have failed or when other forms of manipulation (e.g., asymmetric delay) occur. Thus, detection and defense should be viewed as complementary components of a comprehensive security posture for PTP synchronization.
11. Conclusion
@@
Acknowledgements.
This article is based upon work partially supported by the U.S. National Science Foundation under grants CNS-1925601 and CNS-2112471 and by SERICS (PE00000014) 5GSec project under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. The opinions in the work are solely of the authors and do not reflect those of the the U.S. Military Academy, U.S. Army, or the Department of Defense.Appendix A Appendix
A.1. Additional Attack Results
In this section we present additional results demonstrating the impact of our PTP attacks in terms of reported delay. The delay during normal conditions with FH background traffic is shown in light green, while the delay caused by our attacks is shown in dark red.
Figures 14, 15, and 16 illustrate the impact of our spoofing attack. In Fig.Β 16, the attackβs impact is visible only after it ends because it stops PTP synchronization messages. The longer the attack lasts, the greater the synchronization drift. After the attack, the large spike in delay shows the synchronization drift that occurred.
Figures 17, 18, and 19 show the impact of our replay attack. Unlike the announce attack, PTP traffic continues during the replay attack, allowing delay calculation throughout. However, the reported delay is far higher than the actual network delay, severely impacting PTP synchronization with reported delays over 8000 times greater than the actual observed delay.
A.2. PLFS
In the context of the framework, it is important to discuss PLFS. While FH standardsΒ (OranWG11-secreqspec; etsiFh) allow for any PLFS, in practice, Synchronous Ethernet (SyncE) is commonly used. SyncE is a physical layer protocolΒ (G.8262) that enables high-precision clock synchronization by using the edges in the Ethernet data signal to define the timing content of the signal. This protocol distributes a physical layer clock across a packet network, with each system recovering and forwarding the network timing through the distribution path. SyncE is primarily utilized to assist in frequency synchronization, ensuring that all network elements are aligned to a common frequency reference.
Our production-ready environment employs SyncE, providing robust frequency synchronization to support the timing requirements of PTP. However, despite the presence of SyncE, our attack against PTP was successful, demonstrating the potential vulnerability even with physical layer support. Our digital twin environment does not currently use SyncE, as our focus is on evaluating the security and resilience of the higher layer PTP protocol. This approach allows us to isolate and test the specific impacts of attacks on PTP synchronization without the additional complexity introduced by physical layer frequency support.
A.3. ML Model Parameters
In this section we give a detailed description of each model used.
Logistic Regression:
The Logistic Regression model from scikit-learn is a linear model for binary classification that estimates the probability of a binary outcome using the logistic function. The model is configured with the following parameters: βpenalty=l2β for L2 regularization, βdual=Falseβ to use the primal formulation for the optimization problem, βtol=1e-4β as the tolerance for stopping criteria, βC=1.0β as the inverse of regularization strength, βfit_intercept=Trueβ to add intercepts to the decision function, βsolver=lbfgsβ for optimization, βmax_iter=100β as the maximum number of iterations for the solver to converge.
k-Nearest Neighbors (k-NN):
The k-Nearest Neighbors model is a non-parametric method for classification that identifies the most common class among the βkβ nearest neighbors in the feature space. The model is configured with the following default parameters: βn_neighbors=5β as the number of neighbors to use for k-neighbors queries, βweights=uniformβ to weight all points in each neighborhood equally, βalgorithm=autoβ to automatically select the best algorithm (βball_treeβ, βkd_treeβ, or βbruteβ) based on the input data, βleaf_size=30β as the leaf size for tree-based algorithms, and βp=2β as the power parameter for the Minkowski distance metric, equivalent to the Euclidean distance.
Decision Tree Classifier:
The Decision Tree model is a non-parametric supervised learning method that uses a series of decision rules inferred from data features. The model is configured with βcriterion=giniβ as the splitting criterion based on Gini impurity, βsplitter=bestβ to select the best split, βmax_depth=Noneβ to expand nodes until all leaves are pure or contain fewer than βmin_samples_splitβ samples, βmin_samples_split=2β as the minimum number of samples required to split an internal node, βmin_samples_leaf=1β as the minimum number of samples required at a leaf node, βmax_features=Noneβ to consider all features for the best split, and βrandom_state=Noneβ for randomization.
Gradient Boosting Classifier:
The Gradient Boosting Classifier is an ensemble learning method that builds multiple decision trees sequentially, where each tree corrects the errors of the previous one. The model is configured with βn_estimators=100β as the number of boosting stages (trees), βlearning_rate=0.1β as the step size for weight updates, βmax_depth=3β as the maximum depth of each decision tree, and βverbose=1β to control the verbosity of the training output. Other parameters include βcriterion=friedman_mseβ for measuring the quality of a split, βmin_samples_split=2β for the minimum number of samples required to split an internal node, βmin_samples_leaf=1β for the minimum number of samples required at a leaf node, βmax_features=Noneβ to consider all features for the best split, and βrandom_state=Noneβ for randomization.
Naive Bayes Classifier:
The Gaussian Naive Bayes model is a probabilistic classifier that applies Bayesβ theorem, assuming feature independence given the class. The model is configured with βpriors=Noneβ to use the class priors adjusted according to the data, and βvar_smoothing=1e-9β to set the portion of the largest variance of all features that is added to variances for calculation stability.
Random Forest Classifier:
The Random Forest Classifier is an ensemble learning method that combines multiple decision trees, trained on different parts of the data, for improved accuracy and reduced overfitting. The model is configured with βn_estimators=100β as the number of trees in the forest, βcriterion=giniβ as the splitting criterion based on Gini impurity, βmax_depth=Noneβ to expand nodes until all leaves are pure or contain fewer than βmin_samples_splitβ samples, βmin_samples_split=2β as the minimum number of samples required to split an internal node, βmin_samples_leaf=1β as the minimum number of samples required at a leaf node, βmax_features=sqrtβ setting the number of features to consider when looking for the best split to , and βrandom_state=Noneβ for randomization.
LSTM Classifier:
The LSTM model is designed with a single LSTM layer, followed by a fully connected (FC) layer. The LSTM layer is configured with βbatch_first=Trueβ, allowing the input and output tensors to have the shape β(batch_size, sequence_length, input_dim)β, which facilitates mini-batch training. The LSTM accepts a variable sequence length input ranging from 10 to 40 packets, enabling it to produce a regular, time-based output while processing a variable number of input packets. The hidden layerβs dimensionality is defined by the βhidden_dimβ parameter, and the model outputs a binary classification for each input packet.
The output layer is a fully connected layer that maps the hidden states from the LSTM layer to a single output node for each input packet. The model uses a sigmoid activation function at the final layer to ensure the output is a probability value, making it suitable for binary classification tasks.
During training, the slice length is randomly selected from the range [10, 40] for each epoch. A custom Weighted Binary Cross-Entropy Loss function is employed to handle class imbalance and unequal error costs. For the LSTM model, the loss weight for false positives (FP) is kept constant at 1, while the weight for false negatives (FN) is varied between 1 and 10, achieving maximum recall with a value of 7. The Adam optimizer is used with an initial learning rate of 0.001, and a StepLR scheduler reduces the learning rate by a factor of 0.1 every 10 epochs to help the model converge more effectively. Early stopping is enabled by monitoring validation loss with a patience parameter of 20 epochs.
CNN Model:
The Convolutional Neural Network model is designed with a 2D convolutional architecture. The model processes input sequences of fixed length with features , formatted as a single-channel image of dimensions . The model consists of two convolutional layers.
The first convolutional layer applies 16 filters, each with a kernel size of , using a stride of 1 and padding of 1 to preserve the inputβs spatial dimensions. This is followed by a ReLU activation function and a max pooling operation with a kernel, which reduces the spatial dimensions by a factor of 2. The second convolutional layer uses 32 filters, again with a kernel, and follows the same ReLU activation and max pooling steps. After the second layer, the feature map dimensions are further reduced.
The output from the convolutional layers is flattened to create a vector input for the fully connected (FC) layers. The first FC layer contains 64 neurons with a ReLU activation function, and the final FC layer outputs a single value, which is passed through a sigmoid activation function to produce a binary classification. The model is trained to minimize a custom Weighted Binary Cross-Entropy Loss function to handle class imbalance and uneven costs of false positives and false negatives, although we kept both weights at the default setting of 1.
The Adam optimizer is employed with an initial learning rate of 0.001, and a StepLR scheduler reduces the learning rate by a factor of 0.1 every 10 epochs. Early stopping is enabled by monitoring validation loss with a patience parameter of 20 epochs.
Transformer Model:
The transformer-based model leverages the encoder portion of the transformer architecture, omitting the decoder, to process input sequences with fixed lengths of , , or packets during training and evaluation. These lengths are chosen to balance the trade-off between the need to incorporate more historical data, which improves accuracy and recall, and the associated computational cost and memory requirements.
The encoder is composed of a stack of layers, each layer consisting of a multi-head self-attention mechanism with attention heads, followed by a feed-forward network. The input sequence has six features, and is normalized by a layer normalization step before being processed by the encoder layers. The encoder outputs a contextualized feature matrix of size , where is the number of features and is the sequence length.
The encoded output is then flattened and passed through a fully connected linear layer with 256 neurons, followed by a ReLU activation function to introduce non-linearity. After this, a dropout layer with a dropout probability of 0.2 is applied to mitigate overfitting. The final output is generated by another fully connected layer with a single output neuron, followed by a sigmoid activation function, to produce a binary classification result for the entire input sequence.
We treated the number of attention heads () and the slice length () as hyper-parameters and tuned them for optimal recall. The model is trained using the Adam optimizer with an initial learning rate of 0.001. A StepLR scheduler is used to reduce the learning rate by a factor of 0.1 every 10 epochs, facilitating convergence by fine-tuning the learning rate during training. Early stopping is enabled by monitoring the validation loss with a patience parameter of 20 epochs. We use the same custom Weighted Binary Cross-Entropy Loss function with both weights kept constant at 1.